OpenStudy (anonymous):

what is a simpler form of each product? (4x-6y^3)^2

OpenStudy (anonymous):

\[\left(4 x-6 y^3\right)^2=16 x^2-48 x y^3+36 y^6=4 \left(2 x-3 y^3\right)^2 \]